OpenStudy (anonymous):

@Jadeishere multiply 2(7x+5) a 9x+10 b 9x+7 c 14x+10 d 24x

OpenStudy (anonymous):

C

OpenStudy (anonymous):

ok I have another

OpenStudy (anonymous):

Alright. Are you in Algebra 1?

OpenStudy (anonymous):

no pre al

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

simplify the expression by combining like terms a 2x+6y b 6y c 2x d x+6y

OpenStudy (anonymous):

what is the expression?

OpenStudy (anonymous):

oh sry x-3y+ x +3y

OpenStudy (anonymous):

The -3y cancels out the other 3y. Leaving x+x, so 2x.

OpenStudy (anonymous):

ok I have another

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

simplify the expression by combining like terms -13y-7x-6x

OpenStudy (anonymous):

You cant combine terms that have different variables. So since there is no other Y, you leave the -13y alone. But since both X variables are negative, it would just be adding -7+-6, making the X a -13 too.

OpenStudy (anonymous):

um its minus not negative

OpenStudy (anonymous):

But the minus sign would be the exact same as adding the two numbers with the 6 being negative.

OpenStudy (anonymous):

-7-6= -13 -7+(-6)= -13

OpenStudy (anonymous):

hey here r the answers a -26xy b -13-13x c -13y+1x d -13y-1x

OpenStudy (anonymous):

The only one the makes since is A. None of the variables go away. The answer is -13y+-13x but if you combined that it'd be A. So thats what id choose.

OpenStudy (anonymous):

sense*

OpenStudy (anonymous):

ok thx I have more but I have to go eat ill brb then ok

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

whats up?

OpenStudy (anonymous):

simplify the expression x-23y+x+23y

OpenStudy (anonymous):

2x